OpenStudy (anonymous):

An exam consist of two parts, sec a and sec b. There can be a maximum of 70 questions. There must be at least 10 more questions in sec a that sec b. Write a system of inequalities to model the number of questions in the two sections. let a be the number of questions in sec a and be the number of questions in sec b.

OpenStudy (anonymous):

i need in 4 minutes

OpenStudy (anonymous):

@hsrikako

OpenStudy (anonymous):

a+b = 70 <--Equation a = sec a b= sec b a = b +10 - <--Equation b+(b+10) = 70 2b+10 = 70 2b = 60 b= 30 a+b = 70 a+30 = 70 a=40 b=30

OpenStudy (anonymous):

thanx again